# GenIAS: Generator for Instantiating Anomalies in time Series

### v1

@@ D.1. Notation and Setup
(7) $q_{\phi}(z\mid{\mathbf{X}})=\mathcal{N}\left(\mu,\operatorname{diag}(\sigma^{2})\right),$
@@
(8) $p(z)=\mathcal{N}\left(0,\sigma^{2}_{\text{prior}}I\right),$
where $\sigma^{2}_{\text{prior}}$ is a hyperparameter and $\mu$ is a vector $\mu=(\mu_{1},...,\mu_{L})$. In the following, we consider the effect of using a prior $\sigma^{2}_{\text{prior}}<1$ in our Enhanced KL Loss (Equation 5).
D.2. Prerequisites
Lemma 0 (Compact Latent Space in Enhanced KL Loss).
Consider the KL divergence between the estimated posterior and the prior distributions for a latent dimension $j$. If the enhanced KL loss -Equation 5- is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$. Hence, the variance of latent representations for normal samples are forced to be less than 1, i.e., more “compact”.
@@ D.2. Prerequisites
(9) $D_{\text{KL}}(q(z\mid\mathbf{X})\parallel p(z))=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}+\frac{\mu^{2}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$
@@
(10) $f(\sigma^{2})=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$
@@
(11) $\frac{df}{d\sigma^{2}}=\frac{1}{2}\left(\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}\right).$
@@
(12) $\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}=0\quad\Longrightarrow\quad\sigma^{2}=\sigma^{2}_{\text{prior}}.$
@@
Lemma 0 (Reduced Reconstruction Error for Normal Samples).
@@
(13) $\text{MSE}(\mathbf{X})=\mathbb{E}_{z\sim q_{\phi}(z\mid\mathbf{X})}\left[\|\mathbf{X}-g_{\theta}(z)\|^{2}\right].$
@@
(14) $g_{\theta}(z)\approx g_{\theta}(\mu(\mathbf{X}))+J_{g_{\theta}}(\mu(\mathbf{X}))(z-\mu(\mathbf{X})),$
where $J_{g_{\theta}}$ is the Jacobian of $g_{\theta}$ at $\mu(\mathbf{X})$. We substitute Equation 14 in Equation 13. While $\mu(\mathbf{X})$ stays the same in both Enhanced KL and KL cases, $z-\mu(\mathbf{X})$ varies in these two cases. Hence we focus on the variable term in this substitution, i.e., $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]$. In a multi-dimensional latent space, if the posterior covariance is $\Sigma(\mathbf{X})=\operatorname{diag}(\sigma_{1}^{2}(\mathbf{X}),\dots,\sigma_{d}^{2}(\mathbf{X}))$, then the second-order contribution to the reconstruction error is approximately
(15) $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]=\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right).$
When the enhanced KL loss forces each $\sigma_{j}^{2}(\mathbf{X})$ to be small (namely $\sigma^{2}_{\text{prior}}<1$), the contribution of the variability term
(16) $\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right)$
@@ D.3. Proof of Theorem 1
(17) $q_{\text{normal}}(z)=\mathcal{N}(\mu_{\text{normal}},\operatorname{diag}(\sigma^{2}_{\text{normal}})),$
with $\sigma^{2}_{\text{normal}}\approx\sigma^{2}_{\text{prior}}$ (by Lemma 1). For anomalous inputs, suppose the encoder yields an inflated variance so that
(18) $q_{\text{anom}}(z)=\mathcal{N}(\mu_{\text{anom}},\operatorname{diag}(\psi\sigma^{2}_{\text{normal}})),$
@@
From the given equation the KL divergence between the anomalous and normal posterior distributions in one dimension is:
(19) $\begin{split}D_{\text{KL}}\left(\mathcal{N}(\mu,\psi\sigma^{2}_{\text{normal}})\Big{\|}\mathcal{N}(\mu,\sigma^{2}_{\text{prior}})\right)=\\ \frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}\right).\end{split}$
@@
  • Scenario A (Compact Latent Space): $\sigma_{\text{prior}}<1$, let $\sigma_{\text{prior}}=\alpha$, where $10<\alpha<$ 1.
@@
(20) $D_{\text{KL},A}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right).$
@@
(21) $D_{\text{KL},B}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$
Since $10<\alpha<$ 1, we observe that the fraction $\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}$ in $D_{\text{KL},A}$ is strictly larger than $\frac{\psi\sigma^{2}_{\text{normal}}}{1}$ in $D_{\text{KL},B}$.
@@
(22) $f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right)>f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$
@@
This means that a compact latent space amplifies the separation between normal and anomalous samples under perturbation.
@@ Appendix E Details of Evaluation Metrics
To support a consistent and meaningful evaluation of generated anomalies, we utilize a Deep SVDD (Ruff et al., 2018) representation with a fixed 128-dimensional embedding for all models. This representation provides a structured feature space. Additionally, we evaluate models using test datasets that contain labeled anomalies, allowing for a more reliable assessment of anomaly generation quality. Let $\Omega$ denote the DeepSVDD model. The representation of real anomalous samples in the test dataset is given by $V_{\text{real}}=\Omega(\text{X}_{\text{real anomalies}})$, while the representation of generated anomalous samples, produced by the anomaly generator model from the train dataset, is given by $V_{\text{gen}}=\Omega(\tilde{{\text{X}}})$.
E.1. Anomalous Representation Proximity (ARP)
@@
(24) $\text{ARP}=\frac{1}{1+\frac{1}{|V_{\text{real}}|}\sum_{\mathbf{v}^{r}\in V_{\text{real}}}\min_{\mathbf{v}^{g}\in V_{\text{gen}}}d(\mathbf{v}^{r},\mathbf{v}^{g})}.$
@@
E.2. Entropy-Based Diversity Index (EDI)
@@
(25) $V_{\text{all}}=\bigcup_{m=1}^{M}V_{\text{gen}}^{(m)},$
@@
(26) $p_{i}^{(m)}=\frac{|\{\mathbf{v}_{j}^{(m)}\in V_{\text{gen}}^{(m)}\mid\mathbf{v}_{j}^{(m)}\in R_{i}\}|}{N_{m}},$
where $N_{m}=|V_{\text{gen}}^{(m)}|$ is the total number of generated anomalies for model $m$, and ||$\cdot$|| ⋅ | denotes the cardinality of a set.
@@
E.3. TSAD Evaluation Metrics
We use Best F1 score to measures the performance of TSAD models at the optimal threshold. AUPR and AUROC are holistic measures that assess model performance across various thresholds. All these metrics account for class imbalance, with AUPR placing greater emphasis on anomaly class performance, making it more reflective of anomaly detection capabilities.
@@ Appendix F Implementation Details
Table 6. hyperparameter configurations used in GenIAS.
@@ Appendix G Additional Results
G.1. Visualizing Enhanced KL Loss Effects
@@
Figure 8. Separation of reconstruction error (anomaly score) in the input space with a prior of 0.5 (left) and without a prior (right) on entity C-1 from the MSL dataset. The y-axis is logarithmic.
G.2. TSAD additional results
G.2.1. Critical Difference diagrams
Statistical significance in critical difference digrams was determined using a Friedman test followed by a post-hoc Wilcoxon signed-rank test with Holm correction (Demšar, 2006). The critical difference diagrams in Figure 9 show that GenIAS consistently ranks first across both multivariate (MTS) and univariate (UTS) time series settings for AUPR and F1 Score. In the MTS setting, GenIAS demonstrates a statistically significant improvement over other baselines, indicating its superior performance in capturing anomalies.
Figure 9. Critical difference diagrams for F1 and AUPR scores: (a, b) represent multivariate models, while (c, d) showcase univariate models.
G.2.2. TSAD performance in the other metrics.
@@ G.2.1. Critical Difference diagrams
Table 7. TSAD performance of GenIAS and baseline methods in terms of Aff, PATE and AUROC, with the best model in bold and the second-best model underlined. U and M denote results unavailable for models limited to UTS or MTS, respectively.

### v2

@@ D.1 Notation and Setup
$q_{\phi}(z\mid{\mathbf{X}})=\mathcal{N}\left(\mu,\operatorname{diag}(\sigma^{2})\right),$ (7)
@@
$p(z)=\mathcal{N}\left(0,\sigma^{2}_{\text{prior}}I\right),$ (8)
where $\sigma^{2}_{\text{prior}}$ is a hyperparameter and $\mu$ is a vector $\mu=(\mu_{1},...,\mu_{L})$. In the following, we consider the effect of using a prior $\sigma^{2}_{\text{prior}}<1$ in our Compact KL Loss (Equation 2).
D.2 Prerequisites
Lemma D.1 (Compact Latent Space in Compact KL Loss).
Consider the KL divergence between the estimated posterior and the prior distributions for a latent dimension $j$. If the compact KL loss -Equation 2- is minimized (or its gradient dominates locally) with a prior variance $\sigma^{2}_{\text{prior}}<1$, then for each latent variable $z_{j}$ the optimal posterior variance satisfies $\sigma^{2}_{j}=\sigma^{2}_{\text{prior}}$. Hence, the variance of latent representations for normal samples is forced to be less than 1, i.e., more “compact”.
@@ D.2 Prerequisites
$D_{\text{KL}}(q(z\mid\mathbf{X})\parallel p(z))=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}+\frac{\mu^{2}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$ (9)
@@
$f(\sigma^{2})=\frac{1}{2}\left(\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}-\log\frac{\sigma^{2}}{\sigma^{2}_{\text{prior}}}\right).$ (10)
@@
$\frac{df}{d\sigma^{2}}=\frac{1}{2}\left(\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}\right).$ (11)
@@
$\frac{1}{\sigma^{2}_{\text{prior}}}-\frac{1}{\sigma^{2}}=0\quad\Longrightarrow\quad\sigma^{2}=\sigma^{2}_{\text{prior}}.$ (12)
@@
Lemma D.2 (Reduced Reconstruction Error for Normal Samples).
@@
$\text{MSE}(\mathbf{X})=\mathbb{E}_{z\sim q_{\phi}(z\mid\mathbf{X})}\left[\|\mathbf{X}-g_{\theta}(z)\|^{2}\right].$ (13)
@@
$g_{\theta}(z)\approx g_{\theta}(\mu(\mathbf{X}))+J_{g_{\theta}}(\mu(\mathbf{X}))(z-\mu(\mathbf{X})),$ (14)
where $J_{g_{\theta}}$ is the Jacobian of $g_{\theta}$ at $\mu(\mathbf{X})$. We substitute Equation 14 in Equation 13. While $\mu(\mathbf{X})$ stays the same in both Compact KL and KL cases, $z-\mu(\mathbf{X})$ varies in these two cases. Hence, we focus on the variable term in this substitution, i.e., $\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]$. In a multi-dimensional latent space, if the posterior covariance is $\Sigma(\mathbf{X})=\operatorname{diag}(\sigma_{1}^{2}(\mathbf{X}),\dots,\sigma_{d}^{2}(\mathbf{X}))$, then the second-order contribution to the reconstruction error is approximately
$\mathbb{E}\left[\|J_{g_{\theta}}(z-\mu(\mathbf{X}))\|^{2}\right]=\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right).$ (15)
When the compact KL loss forces each $\sigma_{j}^{2}(\mathbf{X})$ to be small (namely $\sigma^{2}_{\text{prior}}<1$), the contribution of the variability term
$\operatorname{trace}\left(J_{g_{\theta}}^{T}J_{g_{\theta}}\Sigma(\mathbf{X})\right)$ (16)
@@ D.3 Proof of Proposition 1
$q_{\text{normal}}(z)=\mathcal{N}(\mu_{\text{normal}},\operatorname{diag}(\sigma^{2}_{\text{normal}})),$ (17)
with $\sigma^{2}_{\text{normal}}\approx\sigma^{2}_{\text{prior}}$ (by Lemma 3.1). For anomalous inputs, suppose the encoder yields an inflated variance so that
$q_{\text{anom}}(z)=\mathcal{N}(\mu_{\text{anom}},\operatorname{diag}(\psi\sigma^{2}_{\text{normal}})),$ (18)
@@
From the given equation, the KL divergence between the anomalous and normal posterior distributions in one dimension is:
$\begin{split}D_{\text{KL}}\left(\mathcal{N}(\mu,\psi\sigma^{2}_{\text{normal}})\Big\|\mathcal{N}(\mu,\sigma^{2}_{\text{prior}})\right)=\\ \frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\sigma^{2}_{\text{prior}}}\right).\end{split}$ (19)
@@
  • Scenario A (Compact Latent Space): $\sigma_{\text{prior}}<1$, let $\sigma_{\text{prior}}=\alpha$, where $0<\alpha<1$.
@@
$D_{\text{KL},A}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right).$ (20)
@@
$D_{\text{KL},B}=\frac{1}{2}\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}-1-\log\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$ (21)
Since $0<\alpha<1$, we observe that the fraction $\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}$ in $D_{\text{KL},A}$ is strictly larger than $\frac{\psi\sigma^{2}_{\text{normal}}}{1}$ in $D_{\text{KL},B}$.
@@
$f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{\alpha^{2}}\right)>f\left(\frac{\psi\sigma^{2}_{\text{normal}}}{1}\right).$ (22)
@@
This means that a compact latent space amplifies the separation between normal and anomalous samples under perturbation. ∎
@@ Appendix E Details of Evaluation Metrics
[…]
E.1 Anomalous Representation Proximity (ARP)
@@
$\text{ARP}=\frac{1}{1+\frac{1}{|V_{\text{real}}|}\sum_{\mathbf{v}^{r}\in V_{\text{real}}}\min_{\mathbf{v}^{g}\in V_{\text{gen}}}d(\mathbf{v}^{r},\mathbf{v}^{g})},$ (24)
@@
E.2 Entropy-Based Diversity Index (EDI)
@@
$V_{\text{all}}=\bigcup_{m=1}^{M}V_{\text{gen}}^{(m)},$ (25)
@@
$p_{i}^{(m)}=\frac{|\{\mathbf{v}_{j}^{(m)}\in V_{\text{gen}}^{(m)}\mid\mathbf{v}_{j}^{(m)}\in R_{i}\}|}{N_{m}},$ (26)
where $N_{m}=|V_{\text{gen}}^{(m)}|$ is the total number of generated anomalies for model $m$, and $|\cdot|$ denotes the cardinality of a set.
@@
E.3 TSAD Evaluation Metrics
We use the Best F1 score to measure the performance of TSAD models at the optimal threshold. AUPR and AUROC are holistic measures that assess model performance across various thresholds. All these metrics account for class imbalance, with AUPR placing greater emphasis on anomaly class performance, making it more reflective of anomaly detection capabilities.
@@ Appendix F Implementation Details
Table 6: hyperparameter configurations used in GenIAS.
@@ Appendix G Additional Results
G.1 Visualizing Compact KL Loss Effects
@@
Figure 8: Separation of reconstruction error (anomaly score) in the input space with a prior of 0.5 (left) and without a prior (right) on entity C-1 from the MSL dataset. The y-axis is logarithmic.
G.2 TSAD additional results
G.2.1 Critical Difference diagrams
Statistical significance in critical difference diagrams was determined using a Friedman test followed by a post-hoc Wilcoxon signed-rank test with Holm correction (Demšar, 2006). The critical difference diagrams in Figure 9 show that GenIAS consistently ranks first across both multivariate (MTS) and univariate (UTS) time series settings for AUPR and F1 Score. In the MTS setting, GenIAS demonstrates a statistically significant improvement over other baselines, indicating its superior performance in capturing anomalies.
Figure 9: Critical difference diagrams for F1 and AUPR: (a, b) represent multivariate models, while (c, d) showcase univariate models.
G.2.2 TSAD performance in the other metrics.
@@ G.2.1 Critical Difference diagrams
Table 7: TSAD performance of GenIAS and baseline methods in terms of Aff, PATE, and AUROC, with the best model in bold and the second-best model is underlined. U and M denote results unavailable for models limited to UTS or MTS, respectively.